# Follow-Bench: A Unified Motion Planning Benchmark for Socially-Aware Robot Person Following

### v1

@@ II-A Scenarios
We review representative RPF scenarios that pose challenges to both safety and comfort, encompassing both real-world and simulated environments. The difficulty of each scenario is influenced by multiple factors, including the layout complexity of the static environment, the intricacy of the target’s trajectory, and the degree of interference from dynamic objects (i.e., people). Accordingly, we categorize RPF scenarios along three dimensions: target trajectory pattern, dynamic-crowd flow and environmental scale and layout, as shown in Table II-A1.
@@ II-A1 Target Trajectory Pattern
| Targe Trajectory Pattern | Sharp Turn | Triangle trajectory [7, 33] |
@@
| Environmenta Scale & Layout | Scale | Small (10m*10m) [31] |
@@
| Dynamic-Crow Flow | Pedestrian Number | One [38] |
@@ II-B Evaluation Metrics
Comfort is often assessed in terms of motion smoothness, typically evaluated through velocity, acceleration, and jerk [54, 55]. It also depends on the spatial relationship between the robot and the target. Following proxemic theory [19], previous works [33, 7] define a preferred region around the target, as illustrated in Fig. 1. Maintaining a position within the personal zone (0.45-1.2 meters) is generally considered desirable, and researchers often compute the duration the robot remains within this zone. Social awareness further requires maintaining appropriate distances not only from the target but also from surrounding pedestrians. Studies in social navigation [56, 57, 58] suggest maximizing the minimum distance to nearby individuals. For example, Arena [56] introduces a metric that quantifies the time spent within the private zone (less than 0.45 meters) of any pedestrian, with shorter durations indicating better social compliance. Path length is also frequently used as a proxy for energy efficiency [44, 56].
@@ II-C RPF-oriented Motion Planning
Beyond proxemic constraints, collision avoidance is essential. To address this, more sophisticated local planning techniques have been proposed, including dynamic window approach (DWA)[69, 31, 64], SFM[14, 70], model predictive control (MPC)[24, 7, 33, 71], and reinforcement learning (RL)[72, 47, 29, 44]. For example, Dang et al.[64] enhance the DWA by incorporating the desired person-following direction into the velocity sampling process, along with collision constraints and smoothness considerations. SFM methods model the robot’s movement using attractive forces toward the target (or their destination), repulsive forces to avoid obstacles, and social forces to respect interpersonal distance from pedestrians[70].
@@ III-B Scenarios
| Scenario Category | Scenario | Population Levels / Following Distances | Occupancy Variations / Specific Scenarios | ||
| Dynamic-Crowd Flows | Circular crowds | {5, 10, 15, 20, 25, 30} humans | – | ||
| Random crowds | |||||
| Parallel flows | |||||
| Perpendicular flows | |||||
| Environmental Layouts | Corridor | {4, 12, 20, 28} humans (5.6 m width) | Width: {6.8, 6.2, 5.6, 5.0} m (20 humans) | ||
| Intersection | {2, 10, 18, 26} humans (4.8 m width) | Width: {6.0, 5.4, 4.8, 4.2} m (18 humans) | |||
| Doorway | {2, 6, 10, 14} humans (2.8 m width) | Width: {3.4, 3.1, 2.8, 2.5} m (10 humans) | |||
| Cluttered space | {10, 20, 30, 40} humans (30 obstacles) | Obstacles: {40, 30, 20, 10} (30 humans) | |||
| Following Configurations | Default | {Back, Side} (at 1.5 m) | Evaluated in all above scenarios | ||
| Distance variations | {Back, Side} (at {1.0, 1.5, 2.0, 2.5} m) |
In this benchmark, we construct a diverse set of scenarios spanning variations in target trajectories, environmental layouts, and dynamic crowd patterns, drawing upon representative RPF settings summarized in Table II-A1.
@@
To quantify the complexity of socially aware scenarios, we adopt the characterization framework of Stratton et al. [16], evaluating planner performance under varying human densities and environmental occupancy conditions (i.e., passageway widths or obstacle numbers). For dynamic-crowd flows, we define six population levels—5, 10, 15, 20, 25, and 30. For environmental layout scenarios, population levels are set individually according to spatial constraints: Corridor (4, 12, 20, 28) with a width; Intersection (2, 10, 18, 26) with a width; Doorway (2, 6, 10, 14) with a doorway width; and Cluttered space (10, 20, 30, 40) with 30 random obstacles. Furthermore, we vary occupancy conditions by either narrowing passageway widths or increasing obstacle numbers: Corridor (6.8, 6.2, 5.6, 5.0) with 20 humans, Intersection (6.0, 5.4, 4.8, 4.2) with 18 humans, Doorway (3.4, 3.1, 2.8, 2.5) with 10 humans, and Cluttered space (40, 30, 20, 10 obstacles) with 30 humans, respectively. The maximum pedestrian densities, obstacle numbers, and minimum passage widths are determined empirically by observing conditions under which simulated pedestrians become prone to congestion.
In addition, we also evaluate the impact of different person-following positions (back-following and side-following, and different following distances—1.0, 1.5, 2.0, 2.5 meters) on planner performance. These variations aim to assess planners’ adaptability to different social preferences and proxemic requirements. Specifically, we evaluate back- and side-following with following distance in the above all scenarios. While under the variation of following distance, the human number of dynamic-crowd flows is set to 20, and environmental layouts are set to Corridor (20 humans, width), Intersection (18 humans, width), Doorway (10 humans, width), and Cluttered space (30 humans, 30 obstacles). The evaluated scenarios are summarized in Table IV.
@@ III-C Baselines
Proxemic constraints are considered standard practice in recent RPF literature [27, 33, 7]. Accordingly, each planner maintains a “comfortable” following distance of 1.0 m within the target’s personal space. To mitigate occlusions, we adopt adaptive following-position adjustments similar to [7, 74]. In addition, each baseline is equipped with a standard RPF search module [76, 81, 82] that re-finds the target after loss by navigating to the Kalman-filter-predicted position. Given the prominence of MPC-based approaches, we additionally reproduce two MPC variants: (i) MPC w/ trajectory prediction (MPC w/ Traj.) [7], which guides the robot along a predicted target trajectory inferred from historical motion. This planner halts when the predicted trajectory exhibits high uncertainty, thereby prioritizing safety; and (ii) MPC with dynamic search fields (MPC w/ DS.) [53], which employs a dynamic search field to intelligently re-find the target after occlusion—either by overtaking occluders or by temporarily following them until an overtaking opportunity arises.
@@ IV Analysis of RPF planners
Based on the proposed benchmark, we conduct a series of experiments to evaluate the performance of re-implemented RPF planners. The experiments are designed to analyze how these planners perform under all scenarios (Sec. IV-A), different person-following positions, human densities and environmental occupancy conditions (Sec. IV-B). Here, we simulate human-object and human-human interaction based on ORCA. As for experiments with SFM please refer to Appendix. In addition, Appendix also includes complete evaluation results of all experiments and consuming time of each planner.
@@ IV-A Performance under Different Scenarios
We also observe that MPC w/ DS performs comparably to the standalone MPC planner across all scenarios. A noteworthy finding is that their SR and ASR are nearly identical in both the Dynamic and Layout scenarios, indicating that almost all failures arise from collisions, leaving little opportunity for the dynamic search field to take effect. This outcome is expected, as our simulated environments involve complex interactions in which pedestrians’ future trajectories are difficult to predict due to frequent and unpredictable encounters. However, the dynamic search field [53] relies heavily on accurate trajectory prediction of occluders to construct its search field. Consequently, when the predicted field is inaccurate, the resulting search actions may increase the risk of collisions. Meanwhile, the MPC planner [79]—which serves as the backbone of MPC w/ DS.—does not incorporate proactive planning for avoiding dynamic occluders. Thus, once target occlusion occurs, the robot has little room to maneuver for overtaking or following occluders, leaving limited opportunity for the dynamic search field to demonstrate its advantages.
IV-B Performance under Different Following Positions, Human Densities, and Environmental Occupancy Conditions
The results are shown in Fig. 8. In Dynamic and Layout scenarios, back-following consistently outperforms side-following across all planners and metrics. Back-following reduces interference with pedestrian flows, thereby lowering collision risks and improving target visibility. In contrast, side-following is more prone to occlusions and conflicts with pedestrians or static obstacles, leading to reduced SR, ASR, and TVR. Illustrative examples are shown in Fig. 7 (c) and (d), where side-following trajectories significantly interact with obstacles and pedestrian movements, ultimately leading to collisions. However, it is worth noting that side-following is often perceived as more comfortable and socially acceptable in real-world applications, as it allows for better interaction and communication with the target person [46, 20, 61].
@@ VI Conclusions
Robot person following is a fundamental capability in human-robot interaction, with applications spanning personal assistance, security, service, and industrial robotics. This paper provides a comprehensive and structured survey of existing RPF research, with a particular emphasis on two widely recognized and critical criteria: safety and comfort. Guided by these objectives, we review representative scenarios, evaluation metrics, and motion planners. To enable systematic benchmarking of reviewed planners, we introduce Follow-Bench, a unified benchmark implemented in a lightweight simulator that models socially aware interactions, including diverse target trajectory patterns, dynamic-crowd flows, and environmental layouts. To the best of our knowledge, this is the first benchmark to explicitly highlight two essential requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and pedestrians in socially interactive environments. Follow-Bench allows researchers to develop and evaluate new RPF planners, with flexible support for adding maps, obstacles, agents, metrics, and following strategies. All code, datasets, and deployment scripts are publicly released.

### v2

@@ II-A Scenarios
We review representative RPF scenarios that pose challenges to both safety and comfort, encompassing both real-world and simulated environments. The difficulty of each scenario is influenced by multiple factors, including the layout complexity of the static environment, the intricacy of the target’s trajectory, and the degree of interference from dynamic objects (i.e., people). Accordingly, we categorize RPF scenarios along three dimensions: target trajectory pattern, dynamic-crowd flow and environmental scale and layout, as shown in Table I.
@@ II-A1 Target Trajectory Pattern
| Target Trajectory Pattern | Sharp Turn | Triangle trajectory [7, 33] |
@@
| Environmental Scale & Layout | Scale | Small (10m*10m) [31] |
@@
| Dynamic-Crowd Flow | Pedestrian Number | One [38] |
@@ II-B Evaluation Metrics
Comfort is often assessed in terms of motion smoothness, typically evaluated through velocity, acceleration, and jerk [54, 55]. It also depends on the spatial relationship between the robot and the target. Following proxemic theory [19], previous works [33, 7] define a preferred region around the target, as illustrated in Fig. 1. Maintaining a position within the personal zone (0.45-1.2 meters) is generally considered desirable, and researchers often compute the duration the robot remains within this zone. In addition to distance preference, angle preference is also important (e.g., maintaining a side-following position). For instance, Repiso et.al [46] proposes a metric that measures the angular deviation from the preferred following angle.
Social awareness further requires maintaining appropriate distances not only from the target but also from surrounding pedestrians. Studies in social navigation [56, 57, 58] suggest maximizing the minimum distance to nearby individuals. For example, Arena [56] introduces a metric that quantifies the time spent within the private zone (less than 0.45 meters) of any pedestrian, with shorter durations indicating better social compliance. Path length is also frequently used as a proxy for energy efficiency [44, 56].
@@ II-C RPF-oriented Motion Planning
Beyond proxemic constraints, collision avoidance is essential. To address this, more sophisticated local planning techniques have been proposed, including dynamic window approach (DWA)[69, 31, 64], SFM[14, 70], model predictive control (MPC)[24, 7, 33, 71, 72], and reinforcement learning (RL)[73, 47, 29, 44]. For example, Dang et al.[64] enhance the DWA by incorporating the desired person-following direction into the velocity sampling process, along with collision constraints and smoothness considerations. SFM methods model the robot’s movement using attractive forces toward the target (or their destination), repulsive forces to avoid obstacles, and social forces to respect interpersonal distance from pedestrians[70].
@@ III-B Scenarios
| Experiment | Variable Type | Values | Fixed Conditions |
| Target Trajectory | - | - | - |
| Dynamic (H) | Human number | {5, 10, 15, 20, 25, 30} | - |
| Layout (H) | Corridor humans | {4, 12, 20, 28} | 5.6 m width |
| Intersection humans | {2, 10, 18, 26} | 4.8 m width | |
| Doorway humans | {2, 6, 10, 14} | 2.8 m width | |
| Cluttered space humans | {10, 20, 30, 40} | 30 obstacles | |
| Layout (O) | Corridor width | {6.8, 6.2, 5.6, 5.0} m | 20 humans |
| Intersection width | {6.0, 5.4, 4.8, 4.2} m | 18 humans | |
| Doorway width | {3.4, 3.1, 2.8, 2.5} m | 10 humans | |
| Cluttered space obstacles | {10, 20, 30, 40} | 30 humans | |
| Dynamic (F) | Following configuration | Back / Side at {1.0, 1.5, 2.0, 2.5} m | 20 humans |
| Layout (F) | Following configuration | Back / Side at {1.0, 1.5, 2.0, 2.5} m | Corridor: 20 humans; 5.6 m width |
| Intersection: 18 humans; 4.8 m width | |||
| Doorway: 10 humans; 2.8 m width | |||
| Cluttered space: 30 humans; 30 obstacles |
In this benchmark, we construct a diverse set of scenarios spanning variations in target trajectories, environmental layouts, and dynamic crowd patterns, drawing upon representative RPF settings summarized in Table I.
@@
The evaluated scenarios are summarized in Table IV. To quantify the complexity of socially aware scenarios, we adopt the characterization framework of Stratton et al. [16], evaluating planner performance under varying human densities and environmental occupancy conditions (i.e., passageway widths or obstacle numbers). For dynamic-crowd flows, we define six number levels—5, 10, 15, 20, 25, and 30. For environmental layout scenarios, number levels are set individually according to spatial constraints: Corridor (4, 12, 20, 28) with a width; Intersection (2, 10, 18, 26) with a width; Doorway (2, 6, 10, 14) with a doorway width; and Cluttered space (10, 20, 30, 40) with 30 random obstacles. Furthermore, we vary occupancy conditions by either narrowing passageway widths or increasing obstacle numbers: Corridor (6.8, 6.2, 5.6, 5.0) with 20 humans, Intersection (6.0, 5.4, 4.8, 4.2) with 18 humans, Doorway (3.4, 3.1, 2.8, 2.5) with 10 humans, and Cluttered space (40, 30, 20, 10 obstacles) with 30 humans, respectively. The maximum pedestrian densities, obstacle numbers, and minimum passage widths are determined empirically by observing conditions under which simulated pedestrians become prone to congestion.
In addition, we also evaluate the impact of different person-following positions (back-following and side-following, and different following distances—1.0, 1.5, 2.0, 2.5 meters) on planner performance. These variations aim to assess planners’ adaptability to different social preferences and proxemic requirements. Specifically, we evaluate back- and side-following with following distance in the above all scenarios. While under the variation of following distance, the human number of dynamic-crowd flows is set to 20, and environmental layouts are set to Corridor (20 humans, width), Intersection (18 humans, width), Doorway (10 humans, width), and Cluttered space (30 humans, 30 obstacles).
@@ III-C Baselines
Proxemic constraints are considered standard practice in recent RPF literature [27, 33, 7]. Accordingly, each planner maintains a “comfortable” following distance of 1.5 m within the target’s personal space. To mitigate target obstruction, we adopt adaptive following-position adjustments similar to [7, 75]. In addition, each baseline is equipped with a standard RPF search module [77, 82, 83] that re-finds the target after loss by navigating to the Kalman-filter-predicted position. Given the prominence of MPC-based approaches, we additionally reproduce two MPC variants: (i) MPC w/ trajectory prediction (MPC w/ Traj.) [7], which guides the robot along a predicted target trajectory inferred from historical motion. This planner halts when the predicted trajectory exhibits high uncertainty, thereby prioritizing safety; and (ii) MPC with dynamic search fields (MPC w/ DS.) [53], which employs a dynamic search field to intelligently re-find the target after occlusion—either by overtaking occluders or by temporarily following them until an overtaking opportunity arises.
@@ IV Analysis of RPF planners
Based on the proposed benchmark, we conduct a series of experiments to evaluate the performance of re-implemented RPF planners. The experiments are designed to analyze how these planners perform under all scenarios (Sec. IV-A), different person-following distances and positions, human densities and environmental occupancy conditions (Sec. IV-B). Here, we simulate human-object and human-human interaction based on ORCA. For simplicity, we show part of the evaluation metrics here, detailed metric evaluation can be found in the released code. In addition, Appendix includes complete evaluation results of all experiments in each scenario.
@@ IV-A Performance under Different Scenarios
We also observe that MPC w/ DS. performs comparably to the standalone MPC planner across all scenarios. A noteworthy finding is that their SR and ASR are nearly identical in both the Dynamic and Layout scenarios, indicating that almost all failures arise from collisions, leaving little opportunity for the dynamic search field to take effect. This outcome is expected, as our simulated environments involve complex interactions in which pedestrians’ future trajectories are difficult to predict due to frequent and unpredictable encounters. However, the dynamic search field [53] relies heavily on accurate trajectory prediction of occluders to construct its search field. Consequently, when the predicted field is inaccurate, the resulting search actions may increase the risk of collisions. Meanwhile, the MPC planner [80]—which serves as the backbone of MPC w/ DS.—does not incorporate proactive planning for avoiding dynamic occluders. Thus, once target occlusion occurs, the robot has little room to maneuver for overtaking or following occluders, leaving limited opportunity for the dynamic search field to demonstrate its advantages.
IV-B Performance under Different Environmental Settings
The results under different human densities and environmental conditions are shown in Fig. 8. In Dynamic and Layout scenarios, back-following consistently outperforms side-following across all planners and metrics. Back-following reduces interference with pedestrian flows, thereby lowering collision risks and improving target visibility. In contrast, side-following is more prone to occlusions and conflicts with pedestrians or static obstacles, leading to reduced SR, ASR, and TVR. Illustrative examples are shown in Fig. 7 (c) and (d), where side-following trajectories significantly interact with obstacles and pedestrian movements, ultimately leading to collisions. However, it is worth noting that side-following is often perceived as more comfortable and socially acceptable in real-world applications, as it allows for better interaction and communication with the target person [46, 20, 61].
@@ IV-B Performance under Different Environmental Settings
The results under different following distances are presented in Fig. 9. Overall, both overly close (1.0 m) and overly distant (2.0 m or 2.5 m) following distances degrade the performance of the top-performing planners (e.g., MPC-based, MPC w/ Traj., and MPC w/ DS.) across different environmental layouts (Layout (F)). In contrast, this degradation is less pronounced in Dynamic (F); in some cases, the planners even perform better (e.g., MPC-based and MPC w/ DS.) because open spaces provide greater maneuvering room for collision avoidance. The additional room also facilitates person-search, though at farther distances the robot tends to lose track of the target more frequently. Conversely, in Layout (F), narrow environments constrain the available search space, resulting in more collisions. Many of these collisions arise from ambiguities in the preferred following distances relative to environmental structures. For example, in corridors or intersections, some preferred following positions may lie close to walls, thereby creating riskier situations.
Another notable observation from Fig. 9 is that MPC w/ Traj. performs worse than MPC-based and MPC w/ DS. when the following distance is 1.0, 2.0, or 2.5 m. We attribute this to its braking behavior: at short distances (e.g., 1.0 m), abrupt braking increases the likelihood of collisions with surrounding pedestrians. At longer distances (e.g., 2.0 m and 2.5 m), tracking a distant trajectory slows down responsiveness, whereas the point-tracking strategies of MPC-based and MPC w/ DS. allow quicker adjustments. Faster response is also advantageous for re-finding the target after occlusion. Between the two, MPC w/ DS. demonstrates superior person-search performance compared to MPC-based, as the enlarged search space benefits the construction of the dynamic search field, thereby enabling safer and more robust recovery behaviors.
| Metric | MPC | MPC w/ Traj. | MPC w/ DS. | SFM | DWA | DWA w/ Traj. |
| Time (ms) | 55.8 47.7 | 107.2 87.3 | 55.8 47.6 | 0.1 0.0 | 4.3 2.7 | 60.3 35.7 |
IV-C Planning Time Comparison
The computational efficiency results (as shown in Table V) indicate a clear trade-off between planning quality and runtime. The SFM and DWA-based planners exhibit extremely low computation times (below 15 ms), ensuring real-time feasibility. However, their limited optimization capability may compromise planning quality. In contrast, the MPC-based planners require substantially more computation (50-100 ms), with MPC w/ Traj. being the most time-consuming. Although this additional cost may hinder high-frequency control, it potentially provides stronger optimization-based guarantees in complex environments.
IV-D Real-world Experiments
The above experiments indicate that the MPC-based and MPC w/ Traj. planners are the most effective for person-following in complex environments. We implemented these two planners on a differential-drive robot (Scout-mini) equipped with a ZED2 camera for people detection and target identification [6], and a MID-360 LiDAR for obstacle detection and localization using FAST-LIO [84]. The onboard computer is an Intel NUC 11 mini PC with a Core i7-1165G7 CPU. The evaluated scenarios (Fig. 10) include: (i) a target walking randomly with abrupt turns in open space; (ii) a target walking in an environment with static obstacles; and (iii) a target walking in an environment with both static obstacles and moving pedestrians. The preferred following distance was set to 1.5 m, with the robot maintaining a back-following position. Visualizations are provided in the attached video.
In the first scenario, both planners were able to follow the target stably and respond quickly to abrupt maneuvers. However, MPC w/ Traj. produced smoother trajectories and exhibited less aggressive braking due to its trajectory-tracking strategy. In the second scenario, both planners generally avoided collisions with static obstacles, but the MPC-based planner occasionally became stuck when the target executed sharp turns behind obstacles. By contrast, MPC w/ Traj. leveraged predicted trajectories to generate wider-arc motions that avoided close encounters with obstacles. Such wide-arc detours, however, can cause temporary target loss under limited camera FoV and may reduce user comfort. Similar patterns were observed in the third scenario with moving pedestrians. In addition, both planners successfully avoided collisions with sudden pedestrian disturbances by reversing or detouring. These maneuvers, though effective, were reactive rather than proactive since pedestrian trajectory prediction was not incorporated into the planners.
@@ V-A3 Discovering and Utilizing Environmental Patterns
V-A4 Towards More Expressive Intermediate Representations
Intermediate representations continue to be a fundamental challenge in RPF research. Our experiments reveal that point-based or trajectory-based references, coupled with obstacle-avoidance modules, often fail to maintain stable following behavior in crowded environments with frequent occlusions and obstructions. In addition to active following and hierarchical approaches (discussed in Sec. V-A1), an alternative solution is to develop more expressive intermediate representations that integrate multiple information sources and constraints to guide the robot’s motion planning.
A common strategy is to use cost maps [91, 92] or potential fields [93, 37], which provide continuous cost fields that encode environmental obstacles. These methods allow the robot to select follow points or plan paths within feasible regions, rather than relying on isolated instantaneous target points or trajectories. However, they primarily encode obstacle presence and not explicit distances, and updating these maps in real time can be computationally expensive. Euclidean Signed Distance Fields (ESDF) offer a complementary approach by providing a continuous, differentiable distance field that explicitly encodes obstacle distances. ESDFs enable the robot to plan smooth and safe trajectories for efficient target tracking while supporting real-time updates in dynamic environments [94, 95]. However, updating ESDFs in highly dynamic environments can still be computationally demanding.
Beyond collision avoidance, effective intermediate representations must also incorporate additional constraints to improve RPF planning, such as target visibility and social compliance. However, fulfilling all these constraints can result in computational burdens and potentially conflicting objectives. Future research could explore more efficient, adaptive, and expressive representations, for example, by leveraging learning-based methods [96] to encode complex constraints and interactions in a compact form.
@@ VI Conclusions
Robot person following is a fundamental capability in human-robot interaction, with applications spanning personal assistance, security, service, and industrial robotics. This paper provides a comprehensive and structured survey of existing RPF research, with a particular emphasis on two widely recognized and critical criteria: safety and comfort. Guided by these objectives, we review representative scenarios, evaluation metrics, and motion planners. To enable systematic benchmarking of reviewed planners, we introduce Follow-Bench, a unified benchmark implemented in a lightweight simulator that models socially aware interactions, including diverse target trajectory patterns, dynamic-crowd flows, and environmental layouts. To the best of our knowledge, this is the first benchmark to explicitly highlight two essential requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and pedestrians in socially interactive environments. Follow-Bench allows researchers to develop and evaluate new RPF planners, with flexible support for adding maps, obstacles, agents, metrics, and following strategies. All code and deployment scripts are publicly released.
@@
APPENDIX
VI-A Baselines
We re-implement six RPF planners that prioritize both safety and comfort. Specifically, to ensure safety, we incorporate a person-search function in all baselines, which navigates to the Kalman-filter-predicted position of the target when it is lost. To mitigate obstruction, we adopt adaptive following-position adjustments similar to those in [7, 75]. For comfort, we set the default following distance to 1.5 meters, a widely accepted comfortable distance, except for the experiments involving different following configurations.
The SFM-based planner [61] uses the navigation point to estimate the attractive force, while repulsive forces from obstacles and other people are used to compute the resultant force for velocity control. The detailed algorithm is presented in Algorithm 1. More details about force calculation can be found in [61]. The MPC-based planner utilizes the RDA-planner [80] to trace the navigation point while avoiding collisions with static obstacles and pedestrians. We chose the RDA-planner for its efficient collision constraint formulation, which takes the form of a smooth bi-convex dual, facilitating subsequent parallelization. The detailed algorithm is shown in Algorithm 2. The MPC w/ Traj. (outlined in Algorithm 3) extends the MPC-based planner by tracing a reference path based on the target’s predicted future trajectory. Additionally, inspired by [7], we account for the uncertainty in the predicted trajectory to avoid aggressive behaviors, such as the robot braking when the uncertainty exceeds a threshold. Specifically, we employ a Kalman filter to predict the target’s future trajectory and measure the uncertainty using the normalized innovation squared value, which quantifies the difference between the actual observation and the filter’s prediction.
The MPC w/ DS. detailed in Algorithm 4, enhances the MPC-based planner by incorporating a dynamic search field [53], enabling safe person-search after occlusions. The field consists of two modes: overtaking and fluid-following. In overtaking mode, repulsive forces penalize proximity to occluders, while attractive forces encourage minimal travel cost and improved visibility of the target. When overtaking costs exceed a predefined threshold, the planner switches to fluid-following mode, where the crowd is modeled as a flow field. Velocity and density fields are constructed from pedestrian trajectory histories, and the robot’s goal is selected by minimizing a cost function that balances robot velocity, local crowd velocity, and density. This encourages the robot to follow regions of lower density and align with pedestrian flow, thereby improving safety and comfort.
The DWA-based planner [64] samples admissible linear and angular velocities within the dynamic window and performs forward simulation for each candidate. Trajectories are evaluated using a cost function that combines goal alignment, obstacle clearance, and velocity preference, and the optimal velocity is selected as detailed in Algorithm 5. However, this method considers only the robot’s immediate next state, failing to fully exploit the predictive capability of forward simulation and often resulting in shortsighted behaviors. The DWA w/ Traj. (outlined in Algorithm 6) enhances the DWA-based planner by tracing a reference path derived from the target’s predicted future trajectory and refining the cost function. In particular, the heading cost integrates short-term and long-term orientation objectives, while the distance cost ensures real-time target tracking. The trajectory prediction module is identical to that used in the MPC w/ Traj. planner.
VI-B Detailed Experimental Results
Here, we detail our experimental results of every single scenario under different human numbers, occupied conditions and following distances. Fig. 11 and Fig. 12 show the performance under varying human numbers in dynamic crowd flows and environmental layouts, respectively. Fig. 13 shows the performance under different environmental occupied conditions in environmental layouts. Fig. 14 and Fig. 15 show the performance under varying following distances in dynamic crowd flows and environmental layouts, respectively. More evaluation results of other evaluation metrics can be found in the released code.